OpenStudy (anonymous):

rationalize the denominator. Assume that all expressions under raticals represent positive numbers (√(c)-√(d))/(√(c)+√(d))

OpenStudy (callisto):

to rationalize these fraction, just multiply the conjugate (√(c)-√(d))/(√(c)+√(d)) =(√(c)-√(d))/(√(c)+√(d)) x (√(c)-√(d))/(√(c)-√(d)) =(√(c)-√(d))^2 /[(√(c)+√(d)) (√(c)-√(d))] = (√(c)-√(d))^2/(c-d) note that [(√(c)+√(d)) (√(c)-√(d))] can be written as (√(c)^2-√(d)^2) = c-d